# Understanding the Role of Training Data in Test-Time Scaling

### v1

@@ 1 Introduction
Scaling test-time compute enhances inference in large language models (LLMs), by enabling reasoning with long chains-of-thought (CoTs). This allows models to generate more intermediate reasoning steps for complex problems, evaluate multiple options, and backtrack to find more accurate answers, all without changing the model’s parameters. There has been a recent body of work on this idea [15, 20, 11, 22], with OpenAI’s o1 [13] and DeepSeek R1 [6] demonstrating strong reasoning performance with consistent gains from scaling test-time compute. However, our understanding of the training data properties that support test-time scaling remains limited.
Training on diverse and difficult data has shown to be beneficial to enable test-time scaling on complex reasoning tasks, such as mathematical competitions [11], medical reasoning [9], and code [23]. Difficult examples are often identified as those that cannot be answered by the model being trained or other more powerful proxy models. However, the precise notion of difficulty and the relation between the amount of compute at training and test time remains unclear. In particular,
@@
In this paper, we theoretically study the performance of test-time scaling for transformers trained on an in-context weight prediction task for linear regression, where the goal is to predict the linear weight vector from the sequence of input prompts. This framework has been used previously for analyzing the mechanism underlying training CoT [8]. During training, the model performs direct in-context-learning and outputs its prediction of the weight vector. At test time, the transformer performs CoT and generates multiple intermediate steps before arriving at its final prediction of the weight vector. Our analysis yields several intriguing findings: First, fixing the test error, by increasing the test-time compute we can decrease the number of in-context examples (context length) in training prompts. Second, if the skills needed to solve the downstream task (corresponding to directions in the data covariance matrix) are not sufficiently represented in the training data, increasing test-time compute can harm performance, effectively causing the model to overthink. Finally, we characterize hardness of a task based on the smallest eigenvalue of its feature covariance matrix and show that training on a diverse, relevant and hard set of tasks during training yields the best performance for test-time scaling.
@@
    In Section 3, We study in-context learning in transformers with a single linear self-attention (LSA) layer trained via gradient descent. Despite the problem’s non-convexity, we show that gradient descent, when initialized randomly but suitably, converges to a global minimum, which we explicitly characterize. Our analysis allows for general feature covariance. During training, the model engages in direct in-context learning, but at test time we employ chain-of-thought (CoT) prompting to let the model generate intermediate reasoning steps before producing its final output. We demonstrate that, with CoT prompting at test time, the transformer effectively implements a multi-step (pseudo-) Newton’s method for loss optimization. Notably, this part of our contribution extends the results of [24] by incorporating CoT dynamics at test time, and of [8] by accommodating general feature covariance.
@@
    In Section 4, we study a setting with $T$ tasks, where each task is specified by its feature covariance matrix (interpreted, as discussed in part (b), as the set of skills required for the task together with their relative strengths). We extend the analysis of Section 3 to this multi-task setting and characterize the estimation error of the final CoT output. Based on this characterization, we formulate a quadratic optimization problem to determine the optimal task selection probabilities, demonstrating that training on a diverse, relevant, and sufficiently hard set of tasks yields the best performance under test-time scaling. We validate our theoretical results with experiments on both Linear Self-Attention (LSA) models and the more complex nonlinear transformer architecture GPT-2.
@@ 2 Related Work
Recent work has highlighted several phenomena relevant to our study. First, it has been observed that simply increasing test-time compute and reasoning depth can, counterintuitively, harm performance, a phenomenon termed overthinking. The empirical study of [16] suggests that LLMs tend to overthink simple problems by generating unnecessarily long outputs, and underthink harder ones, by providing shallow or incomplete reasoning that overlooks critical steps. In [19], it is argued that exploring more reasoning branches may degrade system efficiency as many branches may be trapped in overthinking. Second, Recent work has explored the test-time scaling paradigm [15, 20], with OpenAI’s o1 [13] and DeepSeek R1 [6] demonstrating strong performance through reinforcement learning on millions of samples and multiple training stages. [11] proposes a simple framework, which involves training on only 1,000 samples with next-token prediction and controlling thinking duration via a simple test-time technique, and show that it achieves test-time scaling and strong reasoning performance. Finally, prior studies on data mixtures emphasize the importance of balancing training corpora with sufficient coverage of topics matched to downstream tasks, as imbalanced data composition can impair generalization [21, 12]. However, prior work has been largely empirical, whereas we develop a theoretical framework that rigorously analyzes test-time scaling and Chain-of-Thought effectiveness, overthinking, and principled strategies for task selection during training.
@@ 3.1 In-Context Weight Prediction and Linear Self-attention
We focus on ICL for linear regression task, where each prompt $P_{\tau}=(x_{\tau,1},y_{\tau,1},\dotsc,x_{\tau,n},y_{\tau,n})$ with $y_{\tau}=\langle w_{\tau},x_{\tau,i}\rangle$, where $x_{\tau,i}\sim{\sf N}(0,\Lambda)$, $w_{\tau}\sim{\sf N}(0,I_{d})$. Most previous works on this setting focus on prediction without directly estimating the weight vector of the test prompt [1, 24, 10]. Here, we take a similar approach to [8] and consider in-context weight prediction where we require the model to directly estimate the wight vector of test prompts. To this end, we adopt the embedding used by [3, 8] which includes the weight-estimation:
@@
where $\psi$ is an activation (e.g. softmax) that is applied column-wise. Following [5, 8, 24, 2], we consider Linear-Self-Attention (LSA) where the activation $\psi$ is the identity mapping. By defining $W:=W_{K}^{\top}W_{Q}$, $V=W_{P}W_{V}$ and $\theta=(W,V)$ we arrive at
@@
Note that chain-of-thought reasoning is not employed during training; however, as we discuss in the following section, the model engages in chain-of-thought reasoning at test time. In contrast, [8, Theorem 3.1] consider the setting of isotropic Gaussian features $(\Lambda=I)$ and incorporate chain-of-thought reasoning during training by generating intermediate steps through gradient updates on the linear regression objective. Also, the result of [24, Theorem 4.1] does not apply to our setting, since it works with a different embedding and trains the model by minimizing the expected prediction loss function.
@@ 3.2 Test time chain-of-thought
$E_{i}=\begin{bmatrix}x_{1}&\cdots&x_{m}&0&0&\dotsc&0\\ y_{1}&\cdots&y_{m}&0&0&\dotsc&0\\ 0&\cdots&0&w_{0}&w_{1}&\dotsc&w_{i}\\ 0&\cdots&0&1&1&\dotsc&1\end{bmatrix}:=\begin{bmatrix}X_{\rm test}&0&0&\dotsc&0\\ y_{\rm test}&0&0&\dotsc&0\\ 0_{d\times n}&w_{0}&w_{1}&\dotsc&w_{i}\\ 0_{1\times n}&1&1&\dotsc&1\end{bmatrix}$ (3.7)
@@ 4 Task selection for training
We next consider a target task with covariance $\Sigma$ (which can be different from any of the tasks during training). For a prompt from it, $P=(x_{1},\langle w_{\rm test},x_{1}\rangle,\dotsc,x_{m},\langle w_{\rm test},x_{m}\rangle)$, we let $X_{\rm test}=[x_{1}|\dotsc|x_{m}]\in\mathbb{R}^{d\times m}$ and $\widehat{\Sigma}:=\frac{1}{m}X_{\rm test}X_{\rm test}^{\top}$. Initializing with $w_{0}=0$ and allowing for a chain-of-thought of length $k$, the LSA estimate of $w_{\rm test}$ reads $w_{k+1}=(I-\Gamma^{-1}\widehat{\Sigma})^{k}w_{\rm test}$. Therefore,
@@ 5 Experiments
Setting. We conduct experiments in two settings. First, we consider a transformer with a single linear self-attention (LSA) to confirm our theory. Then, we consider large, nonlinear transformer architecture namely GPT2 to validate our conclusions. In both cases, the data distribution follows our in-context weight prediction task in Sec. 3.1, where $x_{\tau,i}\sim{\sf N}(0,\Lambda)$, $w_{\tau}\sim{\sf N}(0,I_{d})$. We choose the token dimensions $d=10$. During inference, we let the model to output multiple steps before returning the final predicted weight vector. At each step $i$ we concatenate the embedding with $0_{d},\hat{w}_{i},1]$ as in Eq. (3.7) and input the concatenated embedding matrix to the model. The predicted $w_{k}$ will be returned after $k$ steps of CoT. We report the average results and error bars over 10 runs.
Transformers with a single linear self-attention (LSA). We train the transformer architecture in Eq. (3.2) on the synthetic data generated as described above. We generate 5000 examples, use a batch size $B=1000$ and run Adam with learning rate $\eta=0.001$ for =1000=1000 epochs. For the results reported in the main paper we follow our theoretical setting in Sec. 3.1. That is, we initialize transformer weights as in Eq. (3.6) with $c=1$ and do not perform CoT during training. We report additional results with random initialization and training with CoT in Appendix B.
Large, nonlinear transformer architectures. We use a decoder-only Transformer architecture [17] from the GPT-2 family [14], consisting of 12 layers, 8 attention heads, and a 256-dimensional embedding space. In total the model contains 9.5M parameters.
This architecture takes as input a sequence of vectors in its embedding space and predicts the weight vector within the same space. We apply this architecture to prompts of form $(x_{1},f(x_{1}),\cdots,x_{k},f(x_{k}),w_{0},1)$ in the following manner. In line with [4], we map each $f(x_{i})$ to the same dimension as $x_{i}$ by appending zeros, and map $x_{i},f(x_{i})$ into the latent embedding space of the Transformer through a (learnable) linear transformation. We get the predicted $w_{\tau}$ as the model output. Training is performed with a batch size of 64 over $20k$ total steps. The model is randomly initialized, and CoT is applied during both training and inference.
Larger test-time compute reduces the requirement on training-time compute. Fig 2(a), 2(c) show test error vs length of CoT ($k$). For the LSA model , we use $n=10,20,30$ and for GPT-2 we use $n=20,30,40$. We see that by increasing the test-time compute ($k$), we can decrease the length of prompts $n$ during training to get a similar test error.
@@
Figure 2: More test-time compute reduces training-time requirements for (a) one-layer transformer and (c) GPT-2. However, insufficient task coverage in training data makes longer CoTs harmful for (b) one-layer transformer and (d) GPT-2. For GPT-2, the errorbars are std of 10 runs. For LSA, std is negligible as we start from the fixed initialization in Eq. (3.6).
When more thinking hurts. For training, we sample prompt inputs from ${\sf N}(0,\Lambda)$ where $\Lambda$ is a skewed covariance matrix with eigenbasis chosen uniformly at random and $i$th eigenvalue proportional to $1/i$. For test, we sample prompt inputs from ${\sf N}(0,I_{d})$. We normalize the inputs so that their expected squared norm is equal to that of inputs encountered during training. Fig 2(b), 2(d) show that when some of the directions of the target task are not sufficiently present in the training data, allowing for more thinking during test time would hurt the performance. An interesting observation is that when the model is in overthinking regime (Fig 2(b), 2(d)) larger prompt length $n$ yields a higher test loss, while when the model is not overthinking, larger $n$ reduces the test loss (Fig 2(a), 2(c)).
@@ 5.1 Task selection
We design an experiment to illustrate that our method prioritizes diverse and hard tasks. We consider a multi-task setup with four task types, where each type is defined by two parameters $\alpha$ and $B$. These parameters respectively control the decay rate and the support size of the eigenvalues. Specifically, eigenvalues are proportional to $i^{-\alpha}$ for $i\in[B]$ and zero elsewhere. The positions of nonzero eigenvalues are uniformly shuffled within $[d]$, and the eigenvalues are scaled to have unit sum. Here, $B$ captures task diversity, while $\alpha$ captures the task hardness (with larger $\alpha$ producing smaller nonzero eigenvalues, corresponding to harder tasks according to measure (3.11).)
The four training task types are: Easy-Short $(\alpha=0.2,B=20)$, Hard-Short $(\alpha=0.8,B=20)$, Easy-Long $(\alpha=0.2,B=100)$, and Hard-Long $(\alpha=0.8,B=100)$. The target task is set with $\alpha=0.8$ and $B=d=1000$. We generate 50 tasks of each type by randomizing the eigenbases and the support of eigenvalues. We then solve the quadratic optimization problem (4.5) to obtain task selection probabilities $\pi_{\ell}$, for $\ell=1,\dots,200$. Fig 3(a) displays these probabilities, colored by task type, with solid lines indicating their average per type. As shown, harder and more diverse tasks receive higher selection probabilities, while easier, more concentrated tasks are weighted lower. Fig 3(b) further plots selection probability versus task hardness, confirming that harder tasks are indeed favored, consistent with our theoretical analysis in Section 4.

### v2

@@ 1 Introduction
Scaling test-time compute enhances inference in large language models (LLMs), by enabling reasoning with long chains-of-thought (CoTs). This allows models to generate more intermediate reasoning steps for complex problems, evaluate multiple options, and backtrack to find more accurate answers, all without changing the model’s parameters. There has been a recent body of work on this idea (Snell et al., 2024; Welleck et al., 2024; Muennighoff et al., 2025; Yeo et al., 2025), with OpenAI’s o1 (OpenAI, 2024) and DeepSeek R1 (Guo et al., 2025) demonstrating strong reasoning performance with consistent gains from scaling test-time compute. However, our understanding of the training data properties that support test-time scaling remains limited.
Training on diverse and difficult data has shown to be beneficial to enable test-time scaling on complex reasoning tasks, such as mathematical competitions (Muennighoff et al., 2025), medical reasoning (Huang et al., 2025b), and code (Yu et al., 2025). Difficult examples are often identified as those that cannot be answered by the model being trained or other more powerful proxy models. However, the precise notion of difficulty and the relation between the amount of compute at training and test time remains unclear. In particular,
@@
In this paper, we theoretically study the performance of test-time scaling for transformers trained on an in-context weight prediction task for linear regression, where the goal is to predict the linear weight vector from the sequence of input prompts. This framework has been used previously for analyzing the mechanism underlying training CoT (Huang et al., 2025a). During training, the model performs direct in-context-learning and outputs its prediction of the weight vector. At test time, the transformer performs CoT and generates multiple intermediate steps before arriving at its final prediction of the weight vector. Our analysis yields several intriguing findings: First, fixing the test error, by increasing the test-time compute we can decrease the number of in-context examples (context length) in training prompts. Second, if the skills needed to solve the downstream task (corresponding to directions in the data covariance matrix) are not sufficiently represented in the training data, increasing test-time compute can harm performance, effectively causing the model to overthink. Finally, we characterize hardness of a task based on the smallest eigenvalue of its feature covariance matrix and show that training on a diverse, relevant and hard set of tasks during training yields the best performance for test-time scaling.
@@
    In Section 3, We study in-context learning in transformers with a single linear self-attention (LSA) layer trained via gradient descent. Despite the problem’s non-convexity, we show that gradient descent, when initialized randomly but suitably, converges to a global minimum, which we explicitly characterize. Our analysis allows for general feature covariance. During training, the model engages in direct in-context learning, but at test time we employ chain-of-thought (CoT) prompting to let the model generate intermediate reasoning steps before producing its final output. We demonstrate that, with CoT prompting at test time, the transformer effectively implements a multi-step (pseudo-) Newton’s method for loss optimization. Notably, this part of our contribution extends the results of Zhang et al. (2024) by incorporating CoT dynamics at test time, and of Huang et al. (2025a) by accommodating general feature covariance.
@@
    In Section 5.1, we study a setting with $T$ tasks, where each task is specified by its feature covariance matrix (interpreted, as discussed in part (b), as the set of skills required for the task together with their relative strengths). We extend the analysis of Section 3 to this multi-task setting and characterize the estimation error of the final CoT output. Based on this characterization, we formulate a quadratic optimization problem to determine the optimal task selection probabilities, demonstrating that training on a diverse, relevant, and sufficiently hard set of tasks yields the best performance under test-time scaling. We validate our theoretical results with experiments on both Linear Self-Attention (LSA) models and the more complex nonlinear transformer architecture GPT-2.
@@ 2 Related Work
Recent work has highlighted several phenomena relevant to our study. First, it has been observed that simply increasing test-time compute and reasoning depth can, counterintuitively, harm performance, a phenomenon termed overthinking. The empirical study of Su et al. (2025) suggests that LLMs tend to overthink simple problems by generating unnecessarily long outputs, and underthink harder ones, by providing shallow or incomplete reasoning that overlooks critical steps. In Wang et al. (2025), it is argued that exploring more reasoning branches may degrade system efficiency as many branches may be trapped in overthinking. Second, Recent work has explored the test-time scaling paradigm (Snell et al., 2024; Welleck et al., 2024), with OpenAI’s o1 (OpenAI, 2024) and DeepSeek R1 (Guo et al., 2025) demonstrating strong performance through reinforcement learning on millions of samples and multiple training stages. Muennighoff et al. (2025) proposes a simple framework, which involves training on only 1,000 samples with next-token prediction and controlling thinking duration via a simple test-time technique, and show that it achieves test-time scaling and strong reasoning performance. Finally, prior studies on data mixtures emphasize the importance of balancing training corpora with sufficient coverage of topics matched to downstream tasks, as imbalanced data composition can impair generalization (Xie et al., 2023; Nguyen et al., 2024). However, prior work has been largely empirical, whereas we develop a theoretical framework that rigorously analyzes test-time scaling and Chain-of-Thought effectiveness, overthinking, and principled strategies for task selection during training.
@@ 3.1 In-Context Weight Prediction and Linear Self-attention
We focus on ICL for linear regression task, where each prompt $P_{\tau}=(x_{\tau,1},y_{\tau,1},\dotsc,x_{\tau,n},y_{\tau,n})$ with $y_{\tau}=\langle w_{\tau},x_{\tau,i}\rangle$, where $x_{\tau,i}\sim{\sf N}(0,\Lambda)$, $w_{\tau}\sim{\sf N}(0,I_{d})$. Most previous works on this setting focus on prediction without directly estimating the weight vector of the test prompt (Ahn et al., 2023a; Zhang et al., 2024; Mahankali et al., 2023). Here, we take a similar approach to Huang et al. (2025a) and consider in-context weight prediction where we require the model to directly estimate the wight vector of test prompts. To this end, we adopt the embedding used by Bai et al. (2023); Huang et al. (2025a) which includes the weight-estimation:
@@
where $\psi$ is an activation (e.g. softmax) that is applied column-wise. Following Gatmiry et al. (2024); Huang et al. (2025a); Zhang et al. (2024); Ahn et al. (2023b), we consider Linear-Self-Attention (LSA) where the activation $\psi$ is the identity mapping. By defining $W:=W_{K}^{\top}W_{Q}$, $V=W_{P}W_{V}$ and $\theta=(W,V)$ we arrive at
@@
Note that chain-of-thought reasoning is not employed during training; however, as we discuss in the following section, the model engages in chain-of-thought reasoning at test time. In contrast, (Huang et al., 2025a, Theorem 3.1) consider the setting of isotropic Gaussian features $(\Lambda=I)$ and incorporate chain-of-thought reasoning during training by generating intermediate steps through gradient updates on the linear regression objective. Also, the result of (Zhang et al., 2024, Theorem 4.1) does not apply to our setting, since it works with a different embedding and trains the model by minimizing the expected prediction loss function.
@@ 3.2 Test time chain-of-thought
[…]
@@ 4 Task selection for training
Consider a target task with covariance $\Sigma$ (which can be different from any of the tasks during training). For a prompt from it, $P=(x_{1},\langle w_{\rm test},x_{1}\rangle,\dotsc,x_{m},\langle w_{\rm test},x_{m}\rangle)$, we let $X_{\rm test}=[x_{1}|\dotsc|x_{m}]\in\mathbb{R}^{d\times m}$ and $\widehat{\Sigma}:=\frac{1}{m}X_{\rm test}X_{\rm test}^{\top}$. Initializing with $w_{0}=0$ and allowing a chain-of-thought of length $k$,  (A.9) implies that the LSA estimate of $w_{\rm test}$ is $w_{k+1}=(I-(I-\Gamma^{-1}\widehat{\Sigma})^{k})w_{\rm test}$. Therefore,
@@ 5 Experiments
Setting. We conduct experiments in two settings. First, we consider a transformer with a single linear self-attention (LSA) to confirm our theory. Then, we consider large, nonlinear transformer architecture namely GPT-2 to validate our conclusions. In both cases, the data distribution follows our in-context weight prediction task in Sec. 3.1, where $x_{\tau,i}\sim{\sf N}(0,\Lambda)$, $w_{\tau}\sim{\sf N}(0,I_{d})$. We choose the token dimensions $d=10$. During inference, we let the model to output multiple steps before returning the final predicted weight vector. At each step $i$ we concatenate the embedding with $0_{d},\hat{w}_{i},1]$ as in Eq. (3.7) and input the concatenated embedding matrix to the model. The predicted $w_{k}$ will be returned after $k$ steps of CoT. We report the average results and error bars over 10 runs.
Transformers with a single linear self-attention (LSA). We train the transformer architecture in Eq. (3.2) on the synthetic data generated as described above. We generate 5000 examples, use a batch size $B=1000$ and run Adam with learning rate $\eta=0.001$ for =1000=1000 epochs. For the results reported in the main paper we follow our theoretical setting in Sec. 3.1. That is, we initialize transformer weights $(V(0),W(0))$ according to Theorem 3.1 where $V31(0)$ is set randomly with entries independently and uniformly drawn from $[0,1]$, and we set $c=1$. We do not perform CoT during training.
We report additional results with random initialization and training with CoT in Appendix B.
Large, nonlinear transformer architectures. We use a decoder-only Transformer architecture (Vaswani et al., 2017) from the GPT-2 family (Radford et al., 2019), consisting of 12 layers, 8 attention heads, and a 256-dimensional embedding space. In total the model contains 9.5M parameters.
This architecture takes as input a sequence of vectors in its embedding space and predicts the weight vector within the same space. We apply this architecture to prompts of form $(x_{\tau,1},y_{\tau,1},\cdots,x_{\tau,m},y_{\tau,m},w_{0},1)$ in the following manner. In line with (Garg et al., 2022), we map each $y_{\tau,i}$ to the same dimension as $x_{\tau,i}$ by appending zeros, and map $x_{\tau,i},y_{\tau,i}$ into the latent embedding space of the Transformer through a (learnable) linear transformation. We get the predicted $w_{\tau}$ as the model output. Similarly, we map the model output, i.e., $w_{\tau}$ from the latent embedding space of the Transformer to a $d$-dimensional vector through another (learnable) linear transformation. Training is performed with a batch size of 64 over $25k$ total steps. The model weights are randomly initialized, and CoT is applied during both training and inference. We use curriculum learning (Garg et al., 2022) to speed up training.
Larger test-time compute reduces the requirement on training-time compute. Fig 2(a), 2(c) show the test error vs length of CoT ($k$). For the LSA model, we use $n=10,20,30$ and for GPT-2 we use $n=20,30,40$. We see that by increasing the test-time compute ($k$), we can decrease the length of prompts $n$ during training to get a similar test error.
@@
Figure 2: More test-time compute reduces training-time requirements for (a) one-layer transformer and (c) GPT-2. However, insufficient task coverage in training data makes longer CoTs harmful for (b) one-layer transformer and (d) GPT-2. For GPT-2, the errorbars are std of 10 runs. For LSA, std is negligible as we start from the fixed initialization in Eq. (3.6). Same value for $n$ is used in training and test.
When more thinking hurts. For training, we sample prompt inputs from ${\sf N}(0,\Lambda)$ where $\Lambda$ is a skewed covariance matrix with eigenbasis chosen uniformly at random and $i$-th eigenvalue proportional to $1/i$. For test, we sample prompt inputs from ${\sf N}(0,I_{d})$. We normalize the inputs so that their expected squared norm is equal to that of inputs encountered during training. Fig 2(b), 2(d) show that when some of the directions of the target task are not sufficiently present in the training data, allowing for more thinking during test time would hurt the performance. An interesting observation is that when the model is in overthinking regime (Fig 2(b), 2(d)) larger prompt length $n$ yields a higher test loss, while when the model is not overthinking, larger $n$ reduces the test loss (Fig 2(a), 2(c)).
@@ 5.1 Task selection
We design an experiment to illustrate that our method prioritizes diverse and hard tasks. We consider a multi-task setup with four task types, where each type is defined by two parameters $\alpha$ and $B$. These parameters respectively control the decay rate and the support size of the eigenvalues. Specifically, eigenvalues are proportional to $i^{-\alpha}$ for $i\in[B]$ and zero elsewhere. The positions of nonzero eigenvalues are uniformly shuffled within $[d]$, and the eigenvalues are scaled to have unit sum. Here, $B$ captures task diversity, while $\alpha$ captures the task hardness (with larger $\alpha$ producing smaller nonzero eigenvalues, corresponding to harder tasks according to measure (3.11)).
The four training task types are: Easy-Short $(\alpha=0.2,B=20)$, Hard-Short $(\alpha=0.8,B=20)$, Easy-Long $(\alpha=0.2,B=100)$, and Hard-Long $(\alpha=0.8,B=100)$. The target task is set with $\alpha=0.8$ and $B=d=1000$. We generate 50 tasks of each type by randomizing the eigenbases and the support of eigenvalues. We then solve the quadratic optimization problem (4.5) to obtain task selection probabilities $\pi_{\ell}$, for $\ell=1,\dots,200$. Fig 3(a) displays these probabilities, colored by task type, with solid lines indicating their average per type. As shown, harder and more diverse tasks receive higher selection probabilities, while easier, more concentrated tasks are weighted lower. Fig 3(b) further plots selection probability versus task hardness, confirming that harder tasks are indeed favored, consistent with our theoretical analysis in Section 5.1.
@@
To demonstrate the improvement achieved by our task selection procedure, Appendix B.1 reports simulation results showing test error as a function of the test-time thinking length ($k$) for different task selection strategies. The findings indicate that our optimal task selection procedure effectively prevents overthinking.
5.2 Evaluation on real reasoning benchmarks
Finally, to demonstrate the validity of our conclusions on real reasoning tasks, we train Qwen 2.5-7B-Instruct on the OMEGA dataset (Sun et al., 2025). We chose two tasks from the OMEGA dataset, namely GCD and polynomial root reasoning. These tasks are designed such that training on one does not benefit the performance on the other. We fine-tune the base model (Qwen-Base) with RL separately on the training data of GCD and Poly. We call these models Qwen-GCD and Qwen-Poly. We evaluate both models on the test data of GCD. Table 1 shows that, as expected, for the harder tasks that require longer reasoning, all models have a lower performance. However, we see that while shorter test-time thinking (CoT length less than 1k characters) yields a much better performance (+44.69%) on GCD for Qwen-GCD compared to Qwen-Base, it yields a slightly lower performance on GCD for Qwen-Poly, compared to Qwen-Base (-1.39%). Interestingly, when models reason for longer at test-time (between 1k and 2k characters), Qwen-Poly has a much lower performance (-6.37%) compared to Qwen-Base, while Qwen-GCD outperforms Qwen-Base by 11.2%. This confirms our theoretical results that when training and test data are aligned, more thinking helps. But, insufficient task coverage in training data makes longer test-time compute harmful.
Table 1: Average accuracy on GCD for Qwen2.5-7B Instruct (Base), Base model fine-tuned on CGD (Qwen-GCD) and Base model fine-tuned on Poly (Qwen-Poly). For all the models, the accuracy on examples that require longer CoT is lower (compare the second column to the first column). This confirms that examples that require longer CoT are generally more difficult. The % in () shows the fraction of test data with the corresponding test-time CoT length. Notably, shorter CoTs (0-1k) considerably improves the performance of Qwen-GCD ($75\%$ versus $30.39\%$) and slightly harms the performance of Qwen-Poly ($29\%$ versus $30.39\%$). Longer CoTs improve the performance of Qwen-GCD ($38.4\%$ versus $27.2\%$) and significantly harm the performance of Qwen-Poly ($20.83\%$ versus $27.2\%$).
CoT length [0, 1k) [1k, 2k]
Qwen-Base 30.39% (30% data) 27.2% (70% data)
Qwen-GCD 75% (15% data) 38.4% (85% data)
Qwen-Poly 29% (32% data) 20.83% (68% data)